A flask that weighs 345.8 g is filled with 225.0 mL of carbon tetrachloride. The mass of the flask and carbon tetrachloride is found to be 703.55 g. Calculate the density of carbon tetrachloride.


1
Expert's answer
2021-09-03T01:47:02-0400

mass of carbon tetrachloride =703.55-345.8=357.75g

volume of carbon tetrachloride = 225ml

"density=\frac{mass}{volume}=\frac{357.75g}{225ml}=1.59g/ml"